OpenStudy (anonymous):

PLEASE HELP!!! Which of the following is a polynomial with roots -2, -3i, and 3i ? a.) x3 + 4x2 + 9x + 24 b.) x3 - 4x2 + 9x - 24 c.) x3 + 2x2 + 9x + 18 d.) x3 - 2x2 + 9x - 18

OpenStudy (amistre64):

synthetic division might be useful if you know it

OpenStudy (amistre64):

or if you have a calculator, just plug in those values of x to see if they zero out

OpenStudy (amistre64):

x3 + 4x2 + 9x + 24 0 -2 -4 -10 -2 | 1 2 5 [14] not zero, not a root

OpenStudy (amistre64):

x3 - 4x2 + 9x - 24 0 -2 12 -42 -2 | 1 -6 21 (....) not zero, not a root

OpenStudy (amistre64):

x3 + 2x2 + 9x + 18 0 -2 0 -18 -2 | 1 0 9 [0] is zero, so this ones a possible x^2+9 has no real roots left in it, so its not this one

OpenStudy (amistre64):

x3 - 2x2 + 9x - 18 0 -2 8 -18 -2 | 1 -4 17 hmmm, what am i mistyping

OpenStudy (amistre64):

fine, lets try it this way - (-2-x) (-3-x) (3-x) = x^3 +2x^2 -9x -18 so nonr of them

OpenStudy (anonymous):

(x+2)(x+3)(x-3) x^2+3x+2x+6 (x^2+5x+6)(x-3) x^3+5x^2+6x-3x^2-15x-18 x^3+2x^2-9x-18

OpenStudy (anonymous):

That's weird. Like the poster above said, none of the choices work. Did you mistype it?

OpenStudy (phi):

another way to do this question Which of the following is a polynomial with roots -2, -3, and 3 means x= -2, x= -3 and x= 3 or (x+2)=0, (x+3)= 0, (x-3)=0 multiply together: (x+2)(x+3)(x-3)=0 (x+3)(x-3) = x^2 -9 (x+2)(x^2-9) is x^3 + 2x^2 -9x -18 as noted above, none of your choices match this expression. check for typos, and if you can't find a mistake, ask your teacher about this question.

OpenStudy (anonymous):

@phi look at the question again.. sorry I forgot to put the "i"

OpenStudy (phi):

ok then we start with (x + 3i)(x-3i) = x^2 - 9i^2 = x^2 - 9(-1)= x^2 +9 now do (x^2 +9)(x+2)= x^3 +2x^2 + 9x +18

OpenStudy (phi):

notice I used (x+y)(x-y) = x^2 + y^2 to expand (x + 3i)(x-3i) also, used i*i = -1

OpenStudy (anonymous):

okay?

OpenStudy (phi):

The answer is choice c